OpenStudy (anonymous):

solve AC method 3x³-5x²-2x

OpenStudy (anonymous):

step by step

OpenStudy (amistre64):

factor out an x first off

OpenStudy (anonymous):

u can email to me if easier

OpenStudy (amistre64):

a*c = 6; but there are no factor of 6 that subtract to.....ohyes there are :)

OpenStudy (amistre64):

aint gonna have internet after 7..

OpenStudy (amistre64):

1) after factoring out the "x" we get; 3x^2 -5x -2 2) set it up as tho it CAN factor: (x /3) (x /3) 3) look at last sign...its a (-); we need a (+)(-)..also means we subtract factors to get the middle 4) look at middle sign; its a (-)...the largest # gets the (-) 5) finish setup (x- L/3) (x+S/3) L = large, S = small 6) a*c...3*2 = 6; factors of 6 then... 1,6 ... 6-1 = 5; dont need to go further we got em.. 7) plug into setup (x -6/3) (x+1/3) 8) reduce till you cant reduce no more... 6/3 = 2...stop!! 1/3 = 1/3....stop!! fill it in :) (x-2)(x+1/3) <- if theres a fraction, put the bottom in front like this: (x-2) (3x+1)

OpenStudy (amistre64):

dont forget the "x" you factored to begin with.. x(x-2)(3x+1)

OpenStudy (anonymous):

ty for your help

OpenStudy (anonymous):

im being told by professor that this is not correct we are to rearrange first then factor GCF